OpenStudy (anonymous):

If y varies directly as x and y = −24 when x = 8, then find y when x = 10.

OpenStudy (hba):



OpenStudy (hba):

When you remove the proportionality sign their will be a constant.Let's say k.

OpenStudy (anonymous):

I got -30 as my final answer. Is that correct?

OpenStudy (hba):

Can you explain me how?

OpenStudy (anonymous):

ughhhh... okay. y=kx -24/8=8k/8 k=-3 y=-3x y=-3(10) y=-30

OpenStudy (hba):

Right

OpenStudy (anonymous):

Thanks!!!!

OpenStudy (hba):

No worries c: